OpenStudy (anonymous):

The first term of a geometric series is 3, and the sum of the first term and the second term is 15. What is the sum of the first six terms? A: 1,023 B: 3,906 C: 4,096 D: 11,718

OpenStudy (akashdeepdeb):

Let a = first term r = common ratio a = 6 a + ar = 15 a(1+r) = 15 r = 3/2 So sum of first 6 terms where n=6 is given by: \(S_n = \frac{a(r^n - 1)}{r-1}\) Nwo put in the values, you'll get the answer. :) Understood? :)

OpenStudy (anonymous):

But isn't the first term 3?